# Twinkle Twinkle Little Star, Roman Sees Where You Are: Predicting Exoplanet Transit Yields in the Rosette Nebula with the Nancy Grace Roman Space Telescope

### v1

@@ 3.2 Isochrone Fit
Figure 3: We show our extinction and reddening corrected Gaia sample (circles) in absolute color-magnitude space and the I. Baraffe et al. (2015) evolutionary track for a $0.8\>M_{\odot}$ star (red line) with points at 0.5 Myr, 1 Myr, 2 Myr, 3 Myr, 4 Myr, 5 Myr, 7 Myr, 10 Myr, and 20 Myr (red circles). The purple circles represent stars with interpolated masses above our completeness limit of $0.8\;M_{\odot}$, whereas the gray circles represent members below this limit that we discard in our determination of the initial mass function normalization. We also plot a 10 Myr solar-metallicity PARSEC isochrone (black line) and show the spectral classifications corresponding to each color bin.
@@
As pre-main sequence stars converge onto the zero age main sequence, they exhibit substantial dispersion in their color-magnitude diagram positionality due to effects such as stellar variability, accretion, magnetic activity, and starspots (C. P. M. Bell et al., 2013; G. Somers & M. H. Pinsonneault, 2015). Further, star formation does not happen instantaneously across a SFR, often being triggered sequentially from feedback from massive stars. This results in a genuine age distribution of a few Myr. K. Mužić et al. (2022) demonstrated that this was also true of the Rosette Nebula region (0.5-10 Myr). In addition to a true dispersion in age, some of the scatter can arise from observational uncertainties, such as challenging extinction/reddening constraints and binarity.
@@
To fit an isochrone to our membership sample, we first corrected for dust extinction and reddening. Gaia DR3 provides Apsis-based extinction/reddening estimates for only $\simeq 470$ million of nearly two billion sources, so many stars lack Gaia-provided corrections (O. L. Creevey et al., 2023; A. Vallenari et al., 2023). To maximize the number of usable members in this analysis, we use the median $\mathrm{E(G_{BP}-G_{RP})}$ and $\mathrm{A_{G}}$ derived from a hierarchical binning approach for stars that lack individual Apsis-provided corrections. We first compute the uncorrected stellar color and absolute G-magnitude, dividing these into quantile-based bins. For each missing extinction value, we apply a three-tier strategy: (1) use the median from stars in the same 2D color-magnitude bin, (2) if insufficient data exists, use the median from stars with the same color only, or (3) as a final fallback use the global sample median.
@@
For our analysis, we adopt a single fiducial age of 10 Myr at solar metallicity. We present a PARSEC isochrone (A. Bressan et al., 2012; Y. Chen et al., 2014) with these parameters for our sample in Figure 3. Given that star formation occurs in multiple epochs, choosing the upper limit on the Rosette Nebula’s age allows us to consider more stars that have converged onto the main sequence. We find that the bluer stars ($\mathrm{(G_{BP}-G_{RP})_{0}}\lessapprox 0$) are best fit in color-magnitude space with the 10 Myr isochrone.
@@
We assign masses to each star in our sample by interpolating their extinction-corrected CMD positions onto the 10 Myr PARSEC isochrone. Although the intrinsic stellar mass function rises steeply toward lower masses, the observed number of stars drops below $\sim 0.8\;M_{\odot}$, indicating incompleteness in our Gaia sample in this low-mass regime. We therefore adopt $0.8\>M_{\odot}$ as our completeness limit and restrict our initial mass function (IMF) normalization analysis to the 299 member stars (purple points in Figure 3) brighter and bluer than the I. Baraffe et al. (2015) evolutionary track for a star of this mass at solar-metallicity. Stars fainter/redder than this limit (gray points) are excluded to avoid biases from incompleteness at lower masses, with the near-vertical morphology of the $0.8\;M_{\odot}$ evolutionary track also providing a natural mass-partition.
@@
Stellar multiplicity exhibits a well-established positive correlation with stellar mass (G. Duchêne & A. Kraus, 2013), implying that binarity may bias IMF normalizations. We correct for these effects on our sample by accounting for multiplicity fractions from Table 1 in G. Duchêne & A. Kraus (2013) and assume a flat mass-ratio distribution ($q\sim 0.5$). We adopt this simplification because our mass-completion cuts exclude M dwarfs, where $q\simeq 1$ (T. J. Henry & W.-C. Jao, 2024), whereas most companions for OBA stars remain in our high-mass bin so the detailed mass-ratio distribution has a negligible effect on our final count. Finally, this also prevents overcounting for FGK stars (D. Raghavan et al., 2010). This yields us a multiplicity-corrected sample size of 413 stars.
3.3 Initial Mass Function
@@
  • For $0.8\,M_{\odot}\leq m\leq 20\,M_{\odot}$, we use a Salpeter slope of $\alpha=2.35$ (E. E. Salpeter, 1955). K. Mužić et al. (2019) found that the stellar population in NGC 2244 (the core of the Rosette Nebula) is well-described with a high-mass slope close to the Salpeter value. This mass range corresponds to the portion of our sample that is demonstrably complete, and thus use this bin to find the normalization constant, $k=419$. We note that extremely massive stars ($m\gtrapprox 20\,M_{\odot}$) contribute negligibly to this normalization since the Salpeter slope causes their relative numbers to decline steeply.
  • For $0.01\,M_{\odot}\leq m<0.8\,M_{\odot}$, we adopt $\alpha=1.05$ as measured by K. Mužić et al. (2022), who derive this slope by mapping observed luminosity functions to masses using BT-Settl evolutionary tracks (I. Baraffe et al., 2015), combined with PARSEC isochrones (A. Bressan et al., 2012), and bolometric corrections (M. J. Pecaut & E. E. Mamajek, 2013; P. C. Hewett et al., 2006). This prescription provides a physically motivated, empirically tested description of the IMF down to planetary-mass objects.
@@
4 Determining the Exoplanet Detection Limit
@@
4.1 Synthetic Photometry
Figure 4: We present a 10 Myr brown dwarf M. S. Marley et al. (2021) model spectrum (orange line). We also show the normalized transmission of the Roman broadband F146 filter (black line) and its convolution with the spectrum (orange shaded region), corresponding to its effective stimulus.
@@ 4.1 Synthetic Photometry
  • $0.002\;M_{\odot}\leq m<0.02\;M_{\odot}$ ($2.1\;M_{\rm J}\leq m<21\;M_{\rm J}$): Sonora Bobcat self-luminous sub-stellar atmospheres (M. S. Marley et al., 2021) with age-dependent evolutionary models from D. Saumon & M. S. Marley (2008).
  • $0.02\;M_{\odot}\leq m<0.2\;M_{\odot}$: PHOENIX model atmospheres (P. H. Hauschildt et al., 1997, 1999), with age-dependent evolutionary models from I. Baraffe et al. (2015).
  • $0.2\;M_{\odot}\leq m\leq 0.9\;M_{\odot}$: PARSEC isochrones (A. Bressan et al., 2012; Y. Chen et al., 2014).
For hosts on the low-mass end, with $0.002\;M_{\odot}\leq m<0.2\;M_{\odot}$, we perform a grid-search using the mass-appropriate evolutionary models to find spectra that have the closest match to the corresponding $\log g$ and effective temperature. We renormalize our model flux for each star to the Rosette Nebula distance (1400 pc). Then, we convolve the flux with the F146 and F213 transmission curves (B. Cromey et al., 2023) and derive the host’s AB magnitude (J. B. Oke & J. E. Gunn, 1983). We present the convolution of a model (M. S. Marley et al., 2021) model atmosphere and the F146 bandpass in Figure 4.
@@
Lastly, we employ the Roman Exposure Time Calculator via the Pandeia engine (K. M. Pontoppidan et al., 2016) to estimate S/N for each host from our computed AB magnitudes. We report S/N for a 1 hr long integration, implemented as 229 coadded exposures at an 18.97 s cadence. With readout overheads, the elapsed time slightly exceeds 1 hr. We adopt 1 hr as a representative transit duration to benchmark single-transit detectability, acknowledging that actual durations vary and full campaigns will extend over multiple days.
In Table 2, we present the resulting host AB magnitudes and S/N ratios for FFPs, BDs and stars between $0.002\;M_{\odot}<m<0.9\;M_{\odot}$ (the F146 band saturation limit). The corresponding FFP S/N values indicate that such objects would be detectable down to $\sim 2\>M_{\rm J}$ over the two-week baseline. We note that these calculations do not include the effects from dust extinction. While the survey is sensitive to FFPs, the low S/N ratios per hour (5205–20) make transit detections improbable, except in rare cases of near–equal-size binaries on very short orbits. However, the occurrence rate of such systems is expected to be rare (R. M. Canup & W. R. Ward, 2006). While the survey’s sensitivity limits provide context for the detection threshold of low-mass objects, the primary goal of this study is to assess the detectability of transiting exoplanets around stars. Therefore, our filter selection is informed by the stellar-mass sources with markedly higher S/N values. Between the F146 and F213 filters, the F146 filter is the optimal choice for maximizing sensitivity and yielding the most robust transit detection constraints for sources in the Rosette Nebula.
@@
4.2 Transit Detection
While we perform detailed simulations to predict transit yields in Section 5, we first perform a simple calculation to estimate the exoplanet detection limit in the case of a single transit for the Rosette Nebula. To this aim, we combine our synthetic photometry and S/N calculations with analytic models. Consistent with prior investigations of ultra-short-period planets (R. Sanchis-Ojeda et al., 2014; A. Ofir & S. Dreizler, 2013; A. M. S. Smith et al., 2018), we assume an orbital period of $\sim$8 hours for the transiting companion.
We compute transit durations under the assumption of circular orbits and edge-on inclinations, using the equation given by J. N. Winn (2010):
$t_{\mathrm{dur}}=\frac{P}{\pi}\left[\frac{R_{\mathrm{h}}}{a}\frac{\sqrt{(1+k)^{2}-b^{2}}}{\sin{i}}\right]$ (2)
where $P$ is the orbital period, $R_{\mathrm{h}}$ is the radius of the host, $a$ is the semi-major axis, $k$ is the companion to host radius ratio, $b$ is the orbital impact parameter, and $i$ is the inclination. For each host mass (and corresponding radius), we estimated the median transit duration from the system’s $\sim$ Roche-limit up to 0.2 AU to account for a transitional Roche region, rather than a strict cutoff (Z. M. Leinhardt et al., 2012).
Figure 5: We show our $7\sigma$ detection limit on companion radius as a function of host mass using the black solid line. The green shaded region shows the range of detectable exoplanets around stars whereas the gray shaded shows the range of detectable exoplanets around brown dwarfs. The gray circles show all exoplanets from the NASA Exoplanet archive and the purple markers indicate the planets that are <20<20 Myr old in this parameter space.
@@ 4.2 Transit Detection
We find that the higher S/N achieved in the F146 filter enables the detection of transiting exoplanets orbiting late-type stars, where the planetary occurrence rates are the highest (small planets are $3.5\times$ more frequent around M dwarfs than F or G types; G. D. Mulders et al. 2015), as well as up to solar mass hosts. This reaffirms the Rosette Nebula as a strong candidate for exoplanet discovery. Whereas the TEMPO survey is optimized for very-low mass hosts (FFPs and BDs), the distance to the Rosette Nebula extends the F146 saturation limit to more massive stars, allowing the survey to monitor a broader population of young stars. This is particularly significant because short-period planets are expected to be common in young systems (A. C. Rizzuto et al., 2018; T. J. David et al., 2019), and their inflated radii (J. J. Fortney et al., 2007; J. G. Rogers & J. E. Owen, 2021) further enhance detectability. As shown in Figure 5, only two exoplanets younger than 20 Myr have been identified in this parameter space (T. J. David et al., 2016; T. Holczer et al., 2016). The Rosette Nebula’s young age provides a unique opportunity to extend exoplanet demographics to this scarcely-sampled age regime.
5 Transiting Exoplanet Yields with the Rosette Survey
@@
5.1 Modeling Stellar Variability
@@
Figure 6: We demonstrate that we can detect transits even in the presence of stellar variability, which is expected to be ubiquitous among young stars. Top panel: Light curve simulated in the Roman F146 filter (based on spectrally binned JWST/NIRSpec data) with an injected transit. The light curve is fit with a Gaussian Process + transit model, which accurately recovers the injected transit at the injected location. Bottom panel: Detrended light curve showing the recovered transit detection. This injected transit (915 ppm) lies at our $7\sigma$ detection threshold in the photon-noise-limited case discussed previously, demonstrating that this threshold is an appropriate measure of the transits we can detect even in the presence of stellar variability and instrumental noise.
@@
Young stars are known to exhibit substantial variability (M. G. Barber et al., 2024), generally making transit detection more difficult than for their older main-sequence counterparts. When calculating our transit detection sensitivity, we assumed that our capability would be photon-noise limited. However, stellar variability may also become a limiting factor in detection, as it is particularly pronounced at young stellar ages. We therefore adopt a conservative $7\sigma$ detection threshold—consistent with TEMPO I and other transit surveys (J. M. Jenkins, 2002; W. J. Borucki et al., 2011; C. J. Burke et al., 2015)—which provides both reliability and completeness, particularly in the presence of stellar variability in young stars. Here, we briefly demonstrate that we remain capable of detecting transits that meet our $7\sigma$ detection threshold, even in the presence of stellar variability.
@@
We use archival James Webb Space Telescope (JWST) NIRSpec data of a variable star and construct a light curve by spectrally binning to the F146 Roman filter. This particular source exhibits a much faster rotation rate than expected for young stellar objects, so we stretched the light curve by a factor of ten, resulting in a rotation period of 17.7 hr, typical for stars of this mass and age range (W. Herbst et al., 2007). The resulting simulated data are shown by the light-gray points in Figure 6. The flux level and noise properties of this light curve are consistent with those expected for a 19.32 $\mathrm{mag_{AB}}$ (or 0.09 $M_{\odot}$) star in the Rosette Nebula.
@@
We follow the methodology described in M. A. Limbach et al. (2024), using injection recovery with a Gaussian Process (GP) + transit framework to model variability and quantify detection completeness in time-series photometry. To illustrate that we can detect transits at the $7\sigma$ limit, we inject one of the transits identified in our simulations that was near this detection threshold. This transit has a depth of 915 ppm and a duration of 4.29 hours. In the top panel of Figure 6, we show the light curve with the injected transit, along with the fitted transit and variability model, whereas the bottom panel shows the detrended light curve with the successfully recovered transit. As depicted, our transit-injection recovery pipeline successfully recovers this signal, indicating that even in the presence of realistic stellar variability from space-based photometry, $7\sigma$ detections remain robust. While this demonstration considers only a single-transit event, given our proposed long temporal baseline, real observations will capture multiple transits for most short-period planets. Hence, for these systems, we note that our adopted $7\sigma$ threshold is a conservative lower bound on detectability.
@@
5.2 Creating the Model Population

### v2

@@ III.2 Isochrone Fit
Figure 3: We show our extinction and reddening corrected Gaia sample (circles) in absolute color-magnitude space and the Baraffe et al. (2015) evolutionary track for a $0.8\>M_{\odot}$ star (red line) with points at 0.5 Myr, 1 Myr, 2 Myr, 3 Myr, 4 Myr, 5 Myr, 7 Myr, 10 Myr, and 20 Myr (red circles). The purple circles represent stars with interpolated masses above our completeness limit of $0.8\;M_{\odot}$, whereas the gray circles represent members below this limit that we discard in our determination of the initial mass function normalization. We also plot a 10 Myr solar-metallicity PARSEC isochrone (black line) and show the spectral classifications corresponding to each color bin.
@@
As pre-main sequence stars converge onto the zero age main sequence, they exhibit substantial dispersion in their color-magnitude diagram positionality due to effects such as stellar variability, accretion, magnetic activity, and starspots (Bell et al., 2013; Somers and Pinsonneault, 2015). Further, star formation does not happen instantaneously across a SFR, often being triggered sequentially from feedback from massive stars. This results in a genuine age distribution of a few Myr. Mužić et al. (2022) demonstrated that this was also true of the Rosette Nebula region (0.5-10 Myr). In addition to a true dispersion in age, some of the scatter can arise from observational uncertainties, such as challenging extinction/reddening constraints and binarity.
@@
To fit an isochrone to our membership sample, we first corrected for dust extinction and reddening. Gaia DR3 provides Apsis-based extinction/reddening estimates for only ${\sim}eq470$ million of nearly two billion sources, so many stars lack Gaia-provided corrections (Creevey et al., 2023; Vallenari et al., 2023). To maximize the number of usable members in this analysis, we use the median $\mathrm{E(G_{BP}-G_{RP})}$ and $\mathrm{A_{G}}$ derived from a hierarchical binning approach for stars that lack individual Apsis-provided corrections. We first compute the uncorrected stellar color and absolute G-magnitude, dividing these into quantile-based bins. For each missing extinction value, we apply a three-tier strategy: (1) use the median from stars in the same 2D color-magnitude bin, (2) if insufficient data exists, use the median from stars with the same color only, or (3) as a final fallback use the global sample median.
@@
For our analysis, we adopt a single fiducial age of 10 Myr at solar metallicity. We present a PARSEC isochrone (Bressan et al., 2012; Chen et al., 2014) with these parameters for our sample in Figure 3. Given that star formation occurs in multiple epochs, choosing the upper limit on the Rosette Nebula’s age allows us to consider more stars that have converged onto the main sequence. We find that the bluer stars ($\mathrm{(G_{BP}-G_{RP})_{0}}\lessapprox 0$) are best fit in color-magnitude space with the 10 Myr isochrone.
@@
We assign masses to each star in our sample by interpolating their extinction-corrected CMD positions onto the 10 Myr PARSEC isochrone. Although the intrinsic stellar mass function rises steeply toward lower masses, the observed number of stars drops below ${\sim}0.8\;M_{\odot}$, indicating incompleteness in our Gaia sample in this low-mass regime. We therefore adopt $0.8\>M_{\odot}$ as our completeness limit and restrict our initial mass function (IMF) normalization analysis to the 299 member stars (purple points in Figure 3) brighter and bluer than the Baraffe et al. (2015) evolutionary track for a star of this mass at solar-metallicity. Stars fainter/redder than this limit (gray points) are excluded to avoid biases from incompleteness at lower masses, with the near-vertical morphology of the $0.8\;M_{\odot}$ evolutionary track also providing a natural mass-partition.
@@
Stellar multiplicity exhibits a well-established positive correlation with stellar mass (Duchêne and Kraus, 2013), implying that binarity may bias IMF normalizations. We correct for these effects on our sample by accounting for multiplicity fractions from Table 1 in Duchêne and Kraus (2013) and assume a flat mass-ratio distribution ($q{\sim}0.5$). We adopt this simplification because our mass-completion cuts exclude M dwarfs, where $q\simeq 1$ (Henry and Jao, 2024), whereas most companions for OBA stars remain in our high-mass bin so the detailed mass-ratio distribution has a negligible effect on our final count. Finally, this also prevents overcounting for FGK stars (Raghavan et al., 2010). This yields us a multiplicity-corrected sample size of 413 stars.
III.3 Initial Mass Function
@@
  • For $0.8\,M_{\odot}\leq m\leq 20\,M_{\odot}$, we use a Salpeter slope of $\alpha=2.35$ (Salpeter, 1955). Mužić et al. (2019) found that the stellar population in NGC 2244 (the core of the Rosette Nebula) is well-described with a high-mass slope close to the Salpeter value. This mass range corresponds to the portion of our sample that is demonstrably complete, and thus use this bin to find the normalization constant, $k=419$. We note that extremely massive stars ($m\gtrapprox 20\,M_{\odot}$) contribute negligibly to this normalization since the Salpeter slope causes their relative numbers to decline steeply.
  • For $0.01\,M_{\odot}\leq m<0.8\,M_{\odot}$, we adopt $\alpha=1.05$ as measured by Mužić et al. (2022), who derive this slope by mapping observed luminosity functions to masses using BT-Settl evolutionary tracks (Baraffe et al., 2015), combined with PARSEC isochrones (Bressan et al., 2012), and bolometric corrections (Pecaut and Mamajek, 2013; Hewett et al., 2006). This prescription provides a physically motivated, empirically tested description of the IMF down to planetary-mass objects.
@@
IV Determining the Exoplanet Detection Limit
@@
IV.1 Synthetic Photometry
Figure 4: We present a 10 Myr brown dwarf Marley et al. (2021) model spectrum (orange line). We also show the normalized transmission of the Roman broadband F146 filter (black line) and its convolution with the spectrum (orange shaded region), corresponding to its effective stimulus.
@@ IV.1 Synthetic Photometry
  • $0.002\;M_{\odot}\leq m<0.02\;M_{\odot}$ ($2.1\;M_{\rm J}\leq m<21\;M_{\rm J}$): Sonora Bobcat self-luminous sub-stellar atmospheres (Marley et al., 2021) with age-dependent evolutionary models from Saumon and Marley (2008).
  • $0.02\;M_{\odot}\leq m<0.2\;M_{\odot}$: PHOENIX model atmospheres (Hauschildt et al., 1997, 1999), with age-dependent evolutionary models from Baraffe et al. (2015).
  • $0.2\;M_{\odot}\leq m\leq 0.9\;M_{\odot}$: PARSEC isochrones (Bressan et al., 2012; Chen et al., 2014).
For hosts on the low-mass end, with $0.002\;M_{\odot}\leq m<0.2\;M_{\odot}$, we perform a grid-search using the mass-appropriate evolutionary models to find spectra that have the closest match to the corresponding $\log g$ and effective temperature. We renormalize our model flux for each star to the Rosette Nebula distance (1400 pc). Then, we convolve the flux with the F146 and F213 transmission curves (Cromey et al., 2023) and derive the host’s AB magnitude (Oke and Gunn, 1983). We present the convolution of a model (Marley et al., 2021) model atmosphere and the F146 bandpass in Figure 4.
@@
Lastly, we employ the Roman Exposure Time Calculator via the Pandeia engine (Pontoppidan et al., 2016) to estimate S/N for each host from our computed AB magnitudes. We report S/N for a 1 hr long integration, implemented as 229 coadded exposures at an 18.97 s cadence. With readout overheads, the elapsed time slightly exceeds 1 hr. We adopt 1 hr as a representative transit duration to benchmark single-transit detectability, acknowledging that actual durations vary and full campaigns will extend over multiple days.
In Table 2, we present the resulting host AB magnitudes and S/N ratios for FFPs, BDs and stars between $0.002\;M_{\odot}<m<0.9\;M_{\odot}$ (the F146 band saturation limit). The corresponding FFP S/N values indicate that such objects would be detectable down to ${\sim}2\>M_{\rm J}$ over the two-week baseline. We note that these calculations do not include the effects from dust extinction. While the survey is sensitive to FFPs, the low S/N ratios per hour (5205–20) make transit detections improbable, except in rare cases of near–equal-size binaries on very short orbits. However, the occurrence rate of such systems is expected to be rare (Canup and Ward, 2006). While the survey’s sensitivity limits provide context for the detection threshold of low-mass objects, the primary goal of this study is to assess the detectability of transiting exoplanets around stars. Therefore, our filter selection is informed by the stellar-mass sources with markedly higher S/N values. Between the F146 and F213 filters, the F146 filter is the optimal choice for maximizing sensitivity and yielding the most robust transit detection constraints for sources in the Rosette Nebula.
@@
IV.2 Transit Detection
While we perform detailed simulations to predict transit yields in Section V, we first perform a simple calculation to estimate the exoplanet detection limit in the case of a single transit for the Rosette Nebula. To this aim, we combine our synthetic photometry and S/N calculations with analytic models. Consistent with prior investigations of ultra-short-period planets (Sanchis-Ojeda et al., 2014; Ofir and Dreizler, 2013; Smith et al., 2018), we assume an orbital period of $\sim$}8 hours for the transiting companion.
We compute transit durations under the assumption of circular orbits and edge-on inclinations, using the equation given by Winn (2010):
$t_{\mathrm{dur}}=\frac{P}{\pi}\sin^{-1}\left[\frac{R_{\mathrm{h}}}{a}\frac{\sqrt{(1+k)^{2}-b^{2}}}{\sin{i}}\right]$ (2)
where $P$ is the orbital period, $R_{\mathrm{h}}$ is the radius of the host, $a$ is the semi-major axis, $k$ is the companion to host radius ratio, $b$ is the orbital impact parameter, and $i$ is the inclination. For each host mass (and corresponding radius), we estimated the median transit duration from the system’s $\sim$} Roche-limit up to 0.2 AU to account for a transitional Roche region, rather than a strict cutoff (Leinhardt et al., 2012).
Figure 5: We show our $7\sigma$ detection limit on companion radius as a function of host mass using the black solid line. The green shaded region shows the range of detectable exoplanets around stars whereas the gray shaded shows the range of detectable exoplanets around brown dwarfs. The gray circles show all exoplanets from the NASA Exoplanet archive and the purple markers indicate transiting planets that are <20<20 Myr old in this parameter space.
@@ IV.2 Transit Detection
We find that the higher S/N achieved in the F146 filter enables the detection of transiting exoplanets orbiting late-type stars, where the planetary occurrence rates are the highest (small planets are $3.5\times$ more frequent around M dwarfs than F or G types; Mulders et al. 2015), as well as up to solar mass hosts. This reaffirms the Rosette Nebula as a strong candidate for exoplanet discovery. Whereas the TEMPO survey is optimized for very-low mass hosts (FFPs and BDs), the distance to the Rosette Nebula extends the F146 saturation limit to more massive stars, allowing the survey to monitor a broader population of young stars. This is particularly significant because short-period planets are expected to be common in young systems (Rizzuto et al., 2018; David et al., 2019), and their inflated radii (Fortney et al., 2007; Rogers and Owen, 2021) further enhance detectability. As shown in Figure 5, only three transiting exoplanets younger than 20 Myr have been identified (Barber et al., 2024; David et al., 2016; Mann et al., 2022) around FGKM stars. The Rosette Nebula’s young age provides a unique opportunity to extend exoplanet demographics to this scarcely-sampled age regime.
V Transiting Exoplanet Yields with the Rosette Survey
@@
V.1 Modeling Stellar Variability
@@
Figure 6: We demonstrate that we can detect transits even in the presence of stellar variability, which is expected to be ubiquitous among young stars. Top panel: Light curve simulated in the Roman F146 filter (based on spectrally binned JWST/NIRSpec data for a H = 15.8 mag, T6.5 methane-rich BD; Burgasser et al. 1999) with an injected transit. The light curve is fit with a Gaussian Process + transit model, which accurately recovers the injected transit at the injected location. Bottom panel: Detrended light curve showing the recovered transit detection. This injected transit (915 ppm) lies at our $7\sigma$ detection threshold in the photon-noise-limited case discussed previously, demonstrating that this threshold is an appropriate measure of the transits we can detect even in the presence of stellar variability and instrumental noise.
@@
Young stars are known to exhibit substantial variability (Barber et al., 2024), generally making transit detection more difficult than for their older main-sequence counterparts. When calculating our transit detection sensitivity, we assumed that our capability would be photon-noise limited. However, stellar variability may also become a limiting factor in detection, as it is particularly pronounced at young stellar ages. We therefore adopt a conservative $7\sigma$ detection threshold—consistent with TEMPO I and other transit surveys (Jenkins, 2002; Borucki et al., 2011; Burke et al., 2015)—which provides both reliability and completeness, particularly in the presence of stellar variability in young stars. Here, we briefly demonstrate that we remain capable of detecting transits that meet our $7\sigma$ detection threshold, even in the presence of stellar variability.
@@
We use archival James Webb Space Telescope (JWST) NIRSpec data of a variable star and construct a light curve by spectrally binning to the F146 Roman filter. This particular source exhibits a much faster rotation rate than expected for young stellar objects, so we stretched the light curve by a factor of ten, resulting in a rotation period of 17.7 hr, typical for stars of this mass and age range (Herbst et al., 2007). The resulting simulated data are shown by the light-gray points in Figure 6. The flux level and noise properties of this light curve are consistent with those expected for a 19.32 $\mathrm{mag_{AB}}$ (or 0.09 $M_{\odot}$) star in the Rosette Nebula.
@@
We follow the methodology described in Limbach et al. (2024), using injection recovery with a Gaussian Process (GP) + transit framework to model variability and quantify detection completeness in time-series photometry. To illustrate that we can detect transits at the $7\sigma$ limit, we inject one of the transits identified in our simulations that was near this detection threshold. This transit has a depth of 915 ppm and a duration of 4.29 hours. In the top panel of Figure 6, we show the light curve with the injected transit, along with the fitted transit and variability model, whereas the bottom panel shows the detrended light curve with the successfully recovered transit. As depicted, our transit-injection recovery pipeline successfully recovers this signal, indicating that even in the presence of realistic stellar variability from space-based photometry, $7\sigma$ detections remain robust. While this demonstration considers only a single-transit event, given our proposed long temporal baseline, real observations will capture multiple transits for most short-period planets. Hence, for these systems, we note that our adopted $7\sigma$ threshold is a conservative lower bound on detectability.
@@
Young stars also exhibit prominent variability driven by circumstellar extinction events (dippers) and magnetic activity (flares), but these phenomena can often be distinguished from planetary transits based on geometric constraints and morphological differences.
Dippers, which are the quasi-static or aperiodic dimming caused by occulting dust structures in the inner circumstellar disk (Cody et al., 2014; Cody and Hillenbrand, 2018), typically last 0.5–2 days (as also seen in the $\sim$}10 Myr Upper Scorpius region; Ansdell et al., 2016) and exhibit irregular, asymmetric morphologies with variable depths between events. In contrast, planetary transits result in more symmetric dimming in the stars brightness with durations determined by Equation 2. Combining this with Kepler’s third law and assuming an edge-on, circular orbit with $k\ll 1$, we have a relation between orbital period and transit duration.
Using the stellar mass-radius relation from Table 2, a 0.1 $M_{\odot}$ M dwarf with a 12 hr transit corresponds to an orbital period of ${\sim}450$ days. Hence, dippers, with their characteristic long durations, would require orbital periods of several hundred days. Planets with orbits such as these are temporally separated from our sample of recovered transits (Section V.6), which have durations of a few hours, allowing dippers to be distinguished.
Lastly, flares present distinct geometric and morphological signatures when compared to both dippers and transits. Stellar flares are asymmetric brightness increases with characteristic fast-rise (minutes) and exponential-decay (tens of minutes to hours) timescales (Davenport, 2016). This asymmetric temporal profile can be readily distinguished from the symmetric ingress and egress of planetary transits, where the timescales are determined by the planet’s velocity across the stellar disk. In addition to traditional transit-fitting methods, we can use automated flare identification algorithms such as stella (Feinstein et al., 2020) that are convolutional neural networks trained on these morphological features to robustly identify and mask flare events prior to transit searches.
While our injection-recovery demonstration shows that transits remain detectable in the presence of typical stellar variability, we acknowledge that more extreme variability events could impact our detection completeness. For instance, a large-amplitude flare increasing stellar flux by $\sim$}20% would obscure a simultaneous 0.5% planetary transit. However, recent analyses by Householder et al. (2025) and Wilson et al. (2025) demonstrate that (1) even with more extreme variability amplitudes than our demonstrated case, transit detection algorithms do not produce false positives, and (2) detection thresholds remained close to the photon-noise limit rather than being dominated by stellar activity. This suggests that while extreme variability events may impact our detection limit, the time-averaged detection capability over our 14–30 day baseline approaches the photon-noise-limited case without introducing systematic false positives. Nevertheless, we acknowledge that our predicted yields may be somewhat optimistic if the Rosette Nebula hosts a population of extremely active stars with persistent high-amplitude variability exceeding that observed in WISE J1049-5319 AB and WISE J085510.83-071442.5.
V.2 Creating the Model Population